OpenStudy (anonymous):

solve the equation: 2 x 10^(2-x)=5?

mathslover (mathslover):

\[2* 10^{2-x} = 5\] \[2* [2^{2-x} * 5^{2-x}] = 5\] \[2 * 2^{2-x} * 5^{2-x} = 5\] \[2^{2-x+1} * 5^{2-x}=5\] \[2^{3-x} = 5/5^{2-x}\] \[2^{3-x} = 5^{1-2+x}\] \[2^{3-x} = 5^{x-1}\]

mathslover (mathslover):

wait i am solving it more

OpenStudy (maheshmeghwal9):

Answer must be in Log form^_^

OpenStudy (anonymous):

Yeah the answer is 2-log5/2 i just cant figure out how to get it :(

OpenStudy (unklerhaukus):

\[2 \times 10^{2-x}=5\]\[10^{2-x}=\frac 52\]\[\log_{10}10^{2-x}=\log_{10}\frac52\]\[(2-x)\log_{10}10=\log_{10}5-\log+_{10}2\]\[2-x=\log_{10}5-\log+_{10}2\]\[x=\cdots\]

OpenStudy (unklerhaukus):

* \[2-x=\log_{10}5-\log_{10}2\] \[2-x=\log_{10}\frac52\]

OpenStudy (unklerhaukus):

\[x=\]

OpenStudy (anonymous):

so then -x=-2+log5/2 x=2-log5/2!

OpenStudy (lgbasallote):

lol why did you make it complicated @mathslover :P hahah lol :)

OpenStudy (anonymous):

I know I got lost there they prob did too lol

mathslover (mathslover):

@lgbasallote every thing is complicated in the life lolz actually i am not perfect in Logarithm sorry

OpenStudy (lgbasallote):

your steps are actually interesting..amazing manipulations and mastery of the rules on exponents

mathslover (mathslover):

:) thanks but i dont think that they will help in this question